Describe each of the following graphs in the -plane. Give justification.
A. x = 0
B. y = 0
C. x + y = 0
D. xy = 0

Let me see.
For A, the line x = 0 is the y-axis.
For B, the line y = 0 is the x-axis.
For C, the line x + y = 0 tells me that both x and y have the same number. One is negative and the other positive. It also tells me that x and y can both be zero.
For D, the line xy = 0 tells me that either x is 0 or y is 0.
You say?

Answer by ikleyn(52776)

.
Describe each of the following graphs in the -plane. Give justification.
A. x = 0
B. y = 0
C. x + y = 0
D. xy = 0
Let me see.
For A, the line x = 0 is the y-axis.
For B, the line y = 0 is the x-axis.
For C, the line x + y = 0 tells me that both x and y have the same number. One is negative and the other positive. It also tells me that x and y can both be zero.
For D, the line xy = 0 tells me that either x is 0 or y is 0.
You say?
~~~~~~~~~~~~~~~~~~~~~

(A)  is correct.


(B)  is correct.



(C)  To be more accurate, I would say that for (C), x and y have the same absolute values, but different signs.

     The line x + y = 0  is the same as the line  y = -x.


     Geometrically, it is a straight line with the slope -1 through the origin of the coordinate system
     which goes from upper left to down right, making the angle of 135 = 90 + 45 degrees with
     the positive direction of x-axis.



(D)  The graph of equation  xy = 0  is the union of two straight lines.

     One line is x-axis y= 0.

     Another line is y-axis  x= 0.

     The graph also includes the origin of the coordinate system (0,0), which is the intersection of the two mentioned straight lines.

Solved.

I am glad to see your progress.